OpenStudy (anonymous):

Explain, in complete sentences, how you would use the elimination method to solve the following system of equations. Provide the solution to the system. 6x – 5y = 4 9x + 4y = –17

OpenStudy (anonymous):

write it in matrix form (I'm not an artist so I'll just write it this way): 6,-5,4 9,4,-17 write that matrix in reduced row echelon form: 1,0,-1 0,1,-2 x = -1, y = -2

OpenStudy (anonymous):

6x – 5y = 4 9x + 4y = –17 ------------- 4*(6x – 5y = 4) 5*(9x + 4y = –17) ---------------- 24x – 20y = 16 45x + 20y = –85 --------------(add the above 2 eqns) 69x=-69 x=-1 plug x=-1 in 6x – 5y = 4 6*(-1)-5y=4 -6-5y=4 -5y=10 y=-2 so x=-1 and y=-2